# Embodied LLM Agents Learn to Cooperate in Organized Teams

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as integral tools for reasoning, planning, and decision-making, drawing upon their extensive world knowledge and proficiency in language-related tasks. LLMs thus hold tremendous potential for natural language interaction within multi-agent systems to foster cooperation. However, LLM agents tend to over-report and comply with any instruction, which may result in information redundancy and confusion in multi-agent cooperation. Inspired by human organizations, this paper introduces a framework that imposes prompt-based organization structures on LLM agents to mitigate these problems. Through a series of experiments with embodied LLM agents and human-agent collaboration, our results highlight the impact of designated leadership on team efficiency, shedding light on the leadership qualities displayed by LLM agents and their spontaneous cooperative behaviors. Further, we harness the potential of LLMs to propose enhanced organizational prompts, via a Criticize-Reflect process, resulting in novel organization structures that reduce communication costs and enhance team efficiency.
@@ 1 Introduction
Modern intelligent systems, such as autonomous vehicle networks and swarms of drones, often involve complex decision-making processes where multiple agents or components must collaborate seamlessly to achieve specific objectives Wang et al. (2020); Vinyals et al. (2019); Zhang et al. (2019a); Wang et al. (2021). In these systems, communication among the various agents is pivotal, as it dictates the flow of information, coordination of tasks, and overall system performance Zhang et al. (2019b); Guo et al. (2023); Foerster et al. (2016); Das et al. (2019). Agents in traditional multi-agent systems often have to communicate in pre-specified ways, such as exchanging gradients, sharing data, state observations and actions, etc Kim et al. (2020); Lin et al. (2021); Foerster et al. (2016). The emergence of large language models (LLMs) makes it possible for AI agents to communicate and cooperate using natural language, bringing enormous flexibility and potential for more nuanced and human-understandable interactions Park et al. (2023); Hong et al. (2023); Mandi et al. (2023); Chen et al. (2023a).
@@
Despite the flexibility of LLMs, integrating them into practical multi-agent systems remains a challenge. While LLMs are trained and finetuned for text generation and instruction-following, they are not necessarily tailored to multi-agent cooperation. Modern LLMs are prone to over-reporting and obeying instructions, as a by-product of RLHF finetuning (Bai et al., 2022), and they can ignore critical information (Liu et al., 2023a) or be distracted by irrelevant information (Shi et al., 2023), especially when the context is long (see Figure 1 for examples). While recent studies involving agent-based LLMs have demonstrated they are capable of solving problems through multi-agent collaboration Li et al. (2023b); Zhang et al. (2023c); Mandi et al. (2023), it is worth noting that such collaborations often follow predefined patterns designed using heuristics to channel the behavior of the models productively Li et al. (2023b). Creating systems that support free-flowing interaction between LLMs in a way that could potentially scale to include human collaborators is still an open problem.
@@
This paper investigates the collaborative potential of LLM agents working in teams. Drawing on prior studies in human collaboration from cognitive and economic perspectives, there is potential for organizations to be redesigned to more effectively manage the limited attention span within teams, as suggested by Simon et al. (1971), and mitigate individual limitations and enhance overall team performance, as highlighted by Van Zandt (1999) and Vélez et al. (2023). Specifically, we study two research questions. First, what role do organizational structures play in multi-LLM-agent systems? Second, how can we optimize these organizational structures to support efficient multi-agent coordination? By leveraging AutoGen Wu et al. (2023), a generic multi-agent conversation framework, we develop a framework for studying how to best organize embodied LLM agents to communicate and collaborate in physical/simulated non-text environments Zhang et al. (2023c). Our framework offers the flexibility to prompt and organize LLM agents into various team structures, facilitating versatile inter-agent communication. It also serves as a testbed to empirically evaluate the traditional ideas proposed in the organization theory literature.
@@
Our initial experiments in this setting reveal that uncoordinated LLM agents often send redundant and repetitive messages and interrupt others’ actions, leading to chaos (see Fig. 1 and Appendix D). To remedy these issues, we explore organizational structures that allow multiple LLM agents to collaborate and complete a common task efficiently.
@@
The first organizational structure we explore is a hierarchy, a classic object of study in organizational theory March and Simon (1958); Radner (1993); Chisholm (1992); Bolton and Dewatripont (1994); Garicano (2000); Dodds et al. (2003). With a designated leader, LLM agents work more efficiently and collaboratively. For the example of a three-agent team, imposing a leader improves efficiency by up to 30% with almost no extra communication cost (up to 3%), consistent with findings for human organizations Dodds et al. (2003). This also holds true in five-agent cases. Further, LLM agents demonstrated the potential to elect their own leader and adjust leadership dynamically via communication. With proper organizations, LLM agents exhibit a variety of cooperative behaviors that mimic humans. For example, agents can provide constructive suggestions and seek help from others; they can also execute appropriate interactions for a hierarchy such as reporting back on task progress; see Figures 6 and 7 and Appendix C. We also tested human-agent collaboration, and observe that, unsurprisingly, human leaders are much better at coordinating a team of agents when compared to AI agents.
@@
Figure 1: Example of disorganized communication and interruption, without a designated leader. In a team of three GPT-4 agents, two agents engaged in unnecessary communication and made disordered decisions, causing a delay due to the lack of a predefined organization. We identified many more examples including conflicting messages and repetitive communications, see Appendix D.
@@
In addition to testing existing organizational structures, we explore the use of LLMs to improve the organization prompt. To this end, we develop a Criticize-Reflect framework, adopting a dual LLM architecture, to reflect on the team performance and generate improved and novel organization prompts. Through this iterative process, our LLM agents spontaneously form novel, effective team structures, leading to reduced communication costs and improved efficiency; see Figures 8 and 9.
@@ 2.1 LLM Agents
As powerful LLMs inherit abundant world knowledge and also general reasoning ability, there are increasing efforts to deploy LLMs as the reasoning core for decision-making to build human-like autonomous agents (Sun et al., 2023; Zhu et al., 2023; Hao et al., 2023a). This requires observations of the RL environment to be translated into natural language in a way that is easier for language models to process. The reasoning of the language models also needs to be turned into a viable action for execution. Popular prompting techniques for doing so include ReAct (Yao et al., 2022) and Reflexion (Shinn et al., 2023). Other methods that involve fine-tuning the language models have also been explored (Hao et al., 2023b). In addition, various techniques have been proposed to mitigate the biases and constraints of language models, including chain-of-thought reasoning (Wei et al., 2022), external tools (Shen et al., 2023; Patil et al., 2023), external documents (Wang et al., 2023) and skill libraries (Zhu et al., 2023).
@@ 2.2 Multi-Agent Cooperation
Multi-agent cooperation has been extensively studied for decades under various topics such as communication efficiency, planning, leadership, and team dynamics using different platforms (Lowe et al., 2017; Samvelyan et al., 2019; Resnick et al., 2018; Puig et al., 2021) (see recent surveys for detail (Oroojlooy and Hajinezhad, 2023; Zhang et al., 2021; Gronauer and Diepold, 2022)). Previous works mainly focused on communication through continuous vectors (Das et al., 2019) or discrete symbols (Lowe et al., 2017; Jaques et al., 2019).
Recent work (Xu et al., 2023; Zhang et al., 2023d; Wu et al., 2023; Li et al., 2023a) showed that multiple LLM agents or human-agent teams can improve upon single LLM in solving pure text-based tasks, such as creative writing, reasoning, and code generation. Liu et al. (2023b), Hong et al. (2023) and Zheng et al. (2023) further explored agent selection or role assignment to improve the performance.
LLMs have also been applied to multi-agent cooperation for embodied tasks (Agashe et al., 2023; Mandi et al., 2023; Park et al., 2023; Chen et al., 2023a). Besides, Zhang et al. (2023b) proposed an intention inference framework to enhance the cooperation of LLM agents without explicit communication. Li et al. (2023b) investigated LLM-agents collaboration for Theory of Mind inferences tasks with a broadcast-only communication protocol and homogeneous policies. Zhang et al. (2023c) studied embodied multi-agent cooperation in the two-agent and the one-human-one-agent settings. Chen et al. (2023b) explored different fixed communication structures for multi-LLM-robots. These initial explorations are also limited to fixed team structures and are not optimized for communication efficiency. In contrast, our work explores the impact of deploying and optimizing organizational structures, allowing $\geq 3$ agents in a team, for efficient multi-agent communication and cooperation.
@@ 2.3 Prompt Optimization
Language models are sensitive to prompts. The format of the prompt can have a substantial influence on performance (Gao et al., 2020; Wei et al., 2022; Zhou et al., 2022a; Shi et al., 2023; Zou et al., 2023; Qi et al., 2023). Various research efforts have aimed at improving performance through prompt optimization. Typical approaches include heuristic search using language models’ knowledge (Gao et al., 2020; Shin et al., 2020), first-order methods like soft prompt tuning (Lester et al., 2021), and prefix tuning (Li and Liang, 2021). In this work, we focus on obtaining an interpretable prompt in the form of natural language, drawing on insights from Yang et al. (2023), Zhou et al. (2022b), and Pryzant et al. (2023).
@@ 3.1 Architecture and Multi-Agent Communication
We adopt the embodied LLM-agent architecture proposed by Zhang et al. (2023c) and expand it to enable organized teams of $\geq 3$ agents to communicate, plan, and act in physical/simulated environments. Figure 2 illustrates our architecture. Borrow insights from Zhang et al. (2023c), we adopt four standard modules: Configurator, Perception Module, Memory Module, and Execution Module. They are responsible for configuring the agents, translating environmental observations into text, storing & retrieving historical information, and executing actions, respectively (Fig. 2(a)).
@@
Figure 3: Criticize-Reflect architecture for improving organizational structure. The red agent represents the leader in an organization. The Critic evaluates the trajectories and analyzes the agents’ performance. Together with the external costs from the environment, the Coordinator proposes a new organization prompt to improve the team efficiency.
Previous works focused on two-agent cooperation, in which case the communication can be simply treated as an extra action Mandi et al. (2023); Zhang et al. (2023c). In contrast, we aim to enable three or more agents to work in a team and cooperate through organized communication. Thus we design the architecture with several features that facilitate organized multi-agent communication (Figure 2(b)):
  • We impose an organizational structure for the agent team via prompting, i.e., including a textual description as part of the prompt for each round of communication.
@@ 3.2 Criticize-Reflect Method for Improving Organizational Structure
We leverage powerful LLMs to optimize the organization prompt, borrowing insights from (Yang et al., 2023). To do so, we introduce a dual-LLM framework to allow the multi-LLM-agent system to ponder and improve the organizational structure. Figure 3 illustrates the architecture of our framework. It consists of two LLMs:
  • LLM critic: Inspired by the Actor-Critic method of reinforcement learning Konda and Tsitsiklis (1999), we introduce an LLM critic to evaluate the team’s performance based on verbal feedback. The team critic takes as input the dialogue and action history of one episode. Then, the critic analyzes the input and reasons to extract and summarize the key steps that are believed to influence the performance in the episode. Also, the critic provides a textual evaluation of agents’ behaviors and the ranking of their leadership. Note that the critic in our method is different from Zhang et al. (2023a) where the critic is a centralized controller offering suggestions to each agent. See the prompts in Appendix A.
  • LLM coordinator: The LLM coordinator takes as input the outputs of the LLM critic as well as cost metrics of previous episodes from the environment. It reflects on these data and generates thoughts based on the analysis of the past episodes. The coordinator then generates three distinct new prompts and chooses the best out of them. For more details please refer to Appendix A.
For each new organization prompt, we run for one episode following the new organizational structure and then return the dialogue and action history to the critic. By alternating between criticizing and reflecting on current structure, the framework discovers more effective, novel organizational structures with self-improvement.
@@ 3.3 Environment Setup
We chose VirtualHome-Social Puig et al. (2018, 2021) as the environment and extended it to support multi-LLM-agent communication and interaction. In this environment, agents are humanoid helpers in a virtual home doing housekeeping, where the tasks include Prepare afternoon tea, Wash dishes, Prepare a meal, Put groceries, Set up a dinner table, etc. For instance, in Figure 1, the agents cooperate to prepare afternoon tea by searching for and transporting task-specific items (chocolate, juice, wine, etc.) to a target location (the coffee table). The environment generates symbolic observations of the objects in the home and their relations. Each agent only observes the objects in the open containers located in her room and teammates in the same room, but she can walk to another room to explore. Any agent can communicate with any other agent, not subject to a range limit.
Each episode starts from an initial state where agents are randomly located in the environment and all containers are closed. The episode terminates when the task is fully completed. To evaluate the team’s efficiency we measure the number of time steps taken to task completion, and we report the average number of tokens communicated between agents per step. In our experiment, each run initializes with an independently randomized state to obtain the mean and a confidence interval. We adopt GPT-4, GPT-3.5-turbo (Ouyang et al., 2022), and Llama2-70B (Touvron et al., 2023) as LLMs in our agents. The temperature is set as 0.8, the maximum number of output tokens is 256, and the number of completion choices to generate is 1.

### v2

@@ Abstract
Large Language Models (LLMs) have emerged as integral tools for reasoning, planning, and decision-making, drawing upon their extensive world knowledge and proficiency in language-related tasks. LLMs thus hold tremendous potential for natural language interaction within multi-agent systems to foster cooperation. However, LLM agents tend to over-report and comply with any instruction, which may result in information redundancy and confusion in multi-agent cooperation. Inspired by human organizations, this paper introduces a framework that imposes prompt-based organization structures on LLM agents to mitigate these problems. Through a series of experiments with embodied LLM agents and human-agent collaboration, our results highlight the impact of designated leadership on team efficiency, shedding light on the leadership qualities displayed by LLM agents and their spontaneous cooperative behaviors. Further, we harness the potential of LLMs to propose enhanced organizational prompts, via a Criticize-Reflect process, resulting in novel organization structures that reduce communication costs and enhance team efficiency [^1] Code is available: https://github.com/tobeatraceur/Organized-LLM-Agents. Project website: https://organized-llm-agents.netlify.app/..
@@ 1 Introduction
Modern intelligent systems, such as autonomous vehicle networks and swarms of drones, often involve complex decision-making processes where multiple agents must collaborate seamlessly to achieve specific objectives [57, 55, 66, 58]. In these systems, communication among the various agents is pivotal, as it dictates the flow of information, coordination of tasks, and overall system performance [69, 14, 10, 8]. Agents in traditional multi-agent systems often have to communicate in pre-specified ways, such as exchanging gradients, sharing data, state observations and actions, etc [20, 27, 10]. The emergence of large language models (LLMs) makes it possible for AI agents to communicate and cooperate using natural language, bringing enormous flexibility and potential for more nuanced and human-understandable interactions [36, 17, 32, 5].
@@
Despite the flexibility of LLMs, integrating them into practical multi-agent systems remains a challenge. While LLMs are trained and finetuned for text generation and instruction-following, they are not necessarily tailored to multi-agent cooperation. Modern LLMs are prone to over-reporting and obeying instructions, as a by-product of RLHF finetuning [2], and they can ignore critical information [28] or be distracted by irrelevant information [46], especially when the context is long (see Figure 1 for examples). While recent studies involving agent-based LLMs have demonstrated they are capable of solving problems through multi-agent collaboration [24, 65, 32], it is worth noting that such collaborations often follow predefined patterns designed using heuristics to channel the behavior of the models productively [24]. Creating systems that support free-flowing interaction between LLMs in a way that could potentially scale to include humans is still an open problem.
@@
This paper investigates the collaborative potential of LLM agents working in teams. Drawing on prior studies in human collaboration from cognitive and economic perspectives, there is potential for organizations to be redesigned to more effectively manage the limited attention span within teams, as suggested by Simon et al. [49], and mitigate individual limitations and enhance overall team performance, as highlighted by Van Zandt [53] and Vélez et al. [54]. Specifically, we study two research questions. First, what role do organizational structures play in multi-LLM-agent systems? Second, how can we optimize these organizational structures to support efficient multi-agent coordination? By leveraging AutoGen [60], a generic multi-agent conversation framework, we develop a framework for studying how to best organize embodied LLM agents to communicate and collaborate in physical/simulated non-text environments [65]. Our framework offers the flexibility to prompt and organize LLM agents into various team structures, facilitating versatile inter-agent communication. It also serves as a testbed to empirically evaluate the traditional ideas proposed in the organization theory literature.
@@
Our initial experiments in this setting reveal that uncoordinated LLM agents often send redundant and repetitive messages and interrupt others’ actions, leading to chaos (see Fig. 1 and Appendix E). To remedy these issues, we explore organizational structures, i.e., the dynamics of information exchange, that allow multiple LLM agents to collaborate and complete a common task efficiently.
@@
The first organizational structure we explore is a hierarchy, a classic object of study in organizational theory [33, 42, 7, 3, 12, 9]. With a designated leader, LLM agents work more efficiently and collaboratively. For the example of a three-agent team, imposing a leader improves efficiency by up to 30% with almost no extra communication cost (up to 3%), consistent with findings for human organizations [9]. This also holds true in five-agent cases. Further, LLM agents demonstrated the potential to elect their own leader and adjust leadership dynamically via communication. With proper organizations, LLM agents exhibit a variety of cooperative behaviors that mimic humans. For example, agents can provide constructive suggestions and seek help from others; they can also execute appropriate interactions for a hierarchy such as reporting back on task progress; see Figures 6, 7 and Appendix D. We also tested human-agent collaboration, and observe that, unsurprisingly, human leaders are much better at coordinating a team of agents when compared to AI agents.
@@
Figure 1: Example of disorganized communication and interruption, without a designated leader. In a team of three GPT-4 agents, two agents engaged in unnecessary communication and made disordered decisions, causing a delay due to the lack of a predefined organization. We identified many more examples including conflicting messages and repetitive communications, see Appendix E.
@@
In addition to testing existing organizational structures, we explore the use of LLMs to improve the organizational prompts. To this end, we develop a Criticize-Reflect framework, adopting a dual LLM architecture, to reflect on the team performance and generate improved and novel organizational prompts. Through this iterative process, our LLM agents spontaneously form novel, effective team structures, leading to reduced communication cost and improved efficiency; see Figures 8 and 9.
To summarize, our main contributions are: 1. We design a novel multi-LLM-agent architecture for $\geq 3$ embodied agents, facilitating flexible communication to implement emergent organizational structures. 2. We develop a Criticize-Reflect framework based on LLMs to improve the organizational prompts automatically. 3. Extensive experiments demonstrate that hierarchical organization improves team efficiency, which aligns well with existing literature on human organizations.
@@ 2.1 LLM Agents
As powerful LLMs inherit abundant world knowledge and also general reasoning ability, there are increasing efforts to deploy LLMs as the reasoning core for decision-making to build human-like autonomous agents [50, 74, 15]. This requires observations of the RL environment to be translated into natural language in a way that is easier for LLMs to process. The reasoning of the LLMs also needs to be turned into a viable action for execution. Popular prompting techniques for doing so include ReAct [63] and Reflexion [48]. Other methods that involve fine-tuning the language models have also been explored [16]. In addition, various techniques have been proposed to mitigate the biases and constraints of LLMs, including chain-of-thought reasoning [59], external tools [45, 37], external documents [56] and skill libraries [74].
@@ 2.2 Multi-Agent Cooperation
Multi-agent cooperation has been extensively studied for decades under various topics such as communication efficiency, planning, leadership, and team dynamics using different platforms [30, 44, 43, 40] (see recent surveys for detail [34, 68, 13]). Previous works mainly focused on communication through continuous vectors [8] or discrete symbols [30, 19]. Recent works [61, 67, 60, 23, 18, 26, 51] showed that multiple LLM agents or human-agent teams can improve upon single LLM in solving pure text-based tasks, such as creative writing, reasoning, and code generation. Other works [29, 17, 71] further explored agent selection or role assignment to improve the performance.
LLMs have also been applied to multi-agent cooperation for embodied tasks [1, 32, 36, 5]. Besides, Zhang et al. [64] proposed an intention inference framework to enhance the cooperation of LLM agents without explicit communication. Li et al. [24] investigated LLM-agents collaboration for Theory of Mind inferences tasks with a broadcast-only communication protocol and homogeneous policies. Zhang et al. [65] studied embodied multi-agent cooperation in the two-agent and the one-human-one-agent settings. Chen et al. [6] explored different fixed communication structures for multi-LLM-robots. Zhao et al. [70] and Chen et al. [4] organized the agents by predefined and fixed communication with a virtual manager. These initial explorations are limited to fixed team structures and are not optimized for communication efficiency. In contrast, our work explores the impact of deploying and optimizing organizational structures, allowing $\geq 3$ agents in a team, for efficient multi-agent communication and cooperation.
@@ 2.3 Prompt Optimization
Language models are sensitive to prompts. The format of the prompt can have a substantial influence on performance [11, 59, 72, 46, 75, 41]. Various research efforts have aimed at prompt optimization. Typical approaches include heuristic search using language models’ knowledge [11, 47], first-order methods like soft prompt tuning [22], and prefix tuning [25]. In this work, we focus on obtaining an interpretable prompt in the form of natural language, drawing on insights from Yang et al. [62], Zhou et al. [73], and Pryzant et al. [38].
@@ 3.1 Architecture and Multi-Agent Communication
We adopt the embodied LLM-agent architecture proposed by Zhang et al. [65] and expand it to enable organized teams of $\geq 3$ agents to communicate, plan, and act in physical/simulated environments. Figure 2 illustrates our architecture. Borrowing insights from Zhang et al. [65], we adopt four standard modules: Configurator, Perception Module, Memory Module, and Execution Module. They are responsible for configuring the agents, translating environmental observations into text, storing & retrieving historical information, and executing actions, respectively (Fig. 2(a)).
@@
Previous works focused on two-agent cooperation, in which case the communication can be simply treated as an extra action [32, 65]. In contrast, we aim to enable three or more agents to work in a team and cooperate through emergent organized communication. Thus we design the architecture with several features that facilitate organized multi-agent communication (Figure 2(b)):
  • We disentangle the communication decision-making from the action decision-making by adopting two separate LLMs as Actor and Communicator.
  • We impose an organizational structure for the agent team via prompting, i.e., including a textual description as part of the prompts for both the Actor and Communicator.
@@ 3.2 Criticize-Reflect Method for Improving Organizational Structure
Figure 3: Criticize-Reflect architecture for improving organizational structure. The red agent represents the leader in a hierarchically-organized team. After the team completes one episode, the Critic evaluates the trajectories and analyzes the agents’ performance. Together with the external costs from the environment, the Coordinator proposes a new organizational prompt to improve the team efficiency. The new prompt will be applied to the next episode to continue the iteration.
We leverage powerful LLMs to optimize the organizational prompt, borrowing insights from [62]. To do so, we introduce a dual-LLM framework to allow the multi-LLM-agent system to ponder and improve the organizational structure. Figure 3 illustrates the architecture of our framework. It consists of two LLMs:
  • LLM critic: Inspired by the Actor-Critic method of reinforcement learning [21], we introduce an LLM critic to evaluate the team’s performance based on verbal feedback. The team critic takes as input the dialogue and action history of one episode. Then, the critic analyzes the input and reasons to extract and summarize the key steps that are believed to influence the performance. Also, the critic provides a textual evaluation of agents’ behaviors and the ranking of their leadership. See the prompts in Appendix A and technical details (including the ranking criteria) in Appendix B.1.
  • LLM coordinator: The LLM coordinator takes as input the outputs of the LLM critic as well as cost metrics (time to task completion and communication cost) of previous episodes from the environment. It reflects on these data and generates thoughts based on the analysis of the past episodes and the initial examples. With the reflection of organizational prompts and their performance, the coordinator proposes a new and different organizational prompt for the next episode. Please refer to Appendix A for the prompts and Appendix B.2 for the details of reflection.
For each new organizational prompt, we run for one episode and then return the dialogue and action history to the critic. By criticizing and reflecting on the prompts iteratively, the framework discovers more effective, novel organizational structures with self-improvement.
@@ 3.3 Environment Setup
We chose VirtualHome-Social [39, 40] as the environment and extended it to support multi-LLM-agent communication and interaction. In this environment, agents are humanoid helpers in a virtual home doing housekeeping, where the tasks include Prepare afternoon tea, Wash dishes, Prepare a meal, Put groceries, Set up a dinner table, etc. For instance, in Figure 1, the agents cooperate to prepare afternoon tea by searching for and transporting task-specific items (chocolate, juice, wine, etc.) to a target location (the coffee table). The environment generates symbolic observations of the objects in the home and their relations. Each agent only observes the objects in the open containers located in her room and teammates in the same room, but she can walk to another room to explore. Any agent can communicate with any other agent, not subject to a range limit.
Each episode starts from an initial state where agents are randomly located in the environment and all containers are closed. The episode terminates when the task is fully completed. To evaluate the team’s efficiency we measure the number of time steps taken to task completion, and we report the average number of tokens communicated between agents per step. In our experiment, each run initializes with an independently randomized state to obtain the mean and a confidence interval. We adopt GPT-4, GPT-3.5-turbo [35], and Llama2-70B [52] as LLMs in our agents. The temperature is set as 0.8, the maximum number of output tokens is 256, and the number of completion choices to generate is 1. In practice, we use two Nvidia 80GB A100 GPUs to do the inference on Llama2-70B.